OpenStudy (anonymous):

Write and solve a linear system. You are painting the white lines around the perimeter of a tennis court. You measure and find that the perimeter is 228 feet and the length is 42 feet longer than the width. A: Write a linear system let W be the width and let L be the length of the tennis court. B: Graph the system

OpenStudy (anonymous):

please explain in simple terms!

OpenStudy (jadeishere):

Did i not help well..? Sorry

Mehek (mehek14):

so the formula of finding the perimeter is \(2L+2W\) correct?

OpenStudy (anonymous):

correct...

Mehek (mehek14):

so you have the perimeter is 228 so \(2L+2W=228\) that is your first equation

OpenStudy (anonymous):

ok

Mehek (mehek14):

now they say that the length is 42 ft longer than the width so \(L=42+W\)

OpenStudy (anonymous):

but how do you make the two work together?

Mehek (mehek14):

\(2L+2W=228\\L=42+W\) instead of L in the first equation, you can put 42 + W so \(2L=2(42+W)\\2(42+W)+2W=228\)

OpenStudy (anonymous):

oh! then you distribute and get \[84 + 2w + 2w = 228\] right?

Mehek (mehek14):

right and 2W + 2W = 4W so 84 + 4W = 228 but we don't need to solve like this because your question says to graph it instead

Mehek (mehek14):

so go to http://desmos.com/calculator and put 2x + 2y = 228 and x = 42 + y

OpenStudy (anonymous):

is it supposed to intersect at (78,36)?

Mehek (mehek14):

ye

OpenStudy (anonymous):

so what does that mean? I don't get it. (78,36) represents what?

Mehek (mehek14):

length = 78 width = 36